Notes Ask Your Teacher ถ7.69 points 1 Previous Answers KatzPSE 1 20 p059 The average kinetic eneroy of an argon atom in a spherical container filled with argon gas and with a volume of 1.26 x 102 m2 is 4.36 x 10-21 ). If the pressure inside the container is 1.95 atm, how many moles of argon are in the container? 11. differs from the correct answer by more than 10%. Double check your calculations, moles Need Help? Rasd Submt Answer Save Progress Practice Another Version
Answer #1

Average Kinetic energy

K=(3/2)kBT

4.36*10-21 = (3/2)*(1.38*10-23)*T

T=210.628 K

1 atm =1.01325*105 pa

From Ideal gas law

PV=nRT

=>n=PV/RT =(1.95*1.01325*105)(1.26*10-2)/(8.314)(210.628)

n=1.4217 moles